OpenStudy (anonymous):

what is 4 over 9 plus 4 over 9?

OpenStudy (australopithecus):

4/9 + 4/9

OpenStudy (cwrw238):

if the denominators (bottom of fraction) are equal just add the numerators

OpenStudy (australopithecus):

\[\frac{4}{9} + \frac{4}{9} = \frac{4+4}{9} = \frac{8}{9}\]

OpenStudy (australopithecus):

if the denominators are the same you can add the numerators